# Analysis and Simulation of a Fluid-Heat System in a Thin, Rough Layer in Contact With a Solid Bulk Domain

### v2

@@ 6 Conclusion
We investigated the effective behavior of a nonlinear Stokes-temperature system in a thin layer, where the nonlinearity arises through the convection term in the heat equation and temperature-dependent viscosity. With the concept of two-scale convergence, an effective model was derived where the effective fluid temperature is given by an interface function. We point out that, for the fluid velocity and pressure, we recovered the results of [12] with an additional coupling to the temperature field through viscosity. Multiple simulation results validate the derived model and show that temperature-dependent viscosity also plays a non-negligible role in the limit $\varepsilon\to 0$.
In reality, other parameters (such as conductivity or heat exchange parameter) may also be temperature dependent. Similarly to the present approach, for example, by decoupling the system and showing the existence of solutions with a fixed-point argument, temperature-dependent $\kappa$ and $\alpha$ could be handled as well. Of course, with fitting assumptions for the parameters. Another aspect that could be investigated further is the current $\varepsilon$-scaling of $\kappa$ in the thin layer. Removing this scaling complicates the analysis and may lead to an effective equation without diffusion on the interface $\Sigma$.
Regarding the motivating application of a grinding process, we could derive a justification for the models used in the literature; see [33]. Currently, we are looking at the idealized scenario of periodic roughness. In future research, we plan to extend this model to real grinding wheels and workpiece surfaces. There, we also plan to investigate the effect of including microstructures in the whole process.
@@ Acknowledgements
TF acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project nr. 281474342/GRK2224/2. Additionally, his research was also partially funded by the DFG – project nr. 439916647.

### v3

@@ 6 Conclusion
We investigated the effective behavior of a nonlinear Stokes-temperature system in a thin layer, where the nonlinearity arises through the convection term in the heat equation and temperature-dependent viscosity. With the concept of two-scale convergence, an effective model was derived where the effective fluid temperature is given by an interface function. Multiple simulation results validate the derived model and show that temperature-dependent viscosity also plays a non-negligible role in the limit $\varepsilon\to 0$. The main contribution of this article is the $L^{\infty}$ estimate for the temperature equation in the thin rough layer and the homogenization of a highly non-linear coupled problem.
In reality, other parameters (such as conductivity or heat exchange parameter) may also be temperature dependent. Similarly to the present approach, for example, by decoupling the system and showing the existence of solutions with a fixed-point argument, temperature-dependent $\kappa$ and $\alpha$ could be handled as well. Of course, with fitting assumptions for the parameters. Another aspect that could be investigated further is the current $\varepsilon$-scaling of $\kappa$ in the thin layer. Removing this scaling complicates the analysis and may lead to an effective equation without diffusion on the interface $\Sigma$, as in [33, 4].
Regarding the motivating application of a grinding process, we could derive a justification for the models used in the literature [4], as explained already in the introduction. Currently, we are looking at a strongly idealized scenario of periodic roughness and non-touching roughness with the bottom boundary. In future research, we plan to extend this model to real grinding wheels and workpiece surfaces. Another interesting direction is studying the impact of direct contact between the workpiece and the wheel on the fluid behavior in the contact zone. Additionally, a multiscale approach, also including microscopic material removal, is currently being investigated, studying the effect of microstructures on the whole process.
@@ Acknowledgements
We thank the anonymous reviewers for their constructive and detailed feedback, which helped to improve the article. TF acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project nr. 281474342/GRK2224/2. Additionally, his research was also partially funded by the DFG – project nr. 439916647.